# Every group is represented by an incidence geometry

### v1

@@ Abstract.
We investigate the relationship between finite groups and incidence geometries through their automorphism structures. Building upon classical results on the realizability of groups as automorphism groups of graphs, we develop a general framework to represent pairs of finite groups $(G,H)$, where $H\trianglelefteq G$, as pairs of correlation–automorphism groups of suitable incidence geometries. Specifically, we prove that for every such pair $(G,H)$, there exists an incidence geometry $\Gamma$ satisfying that the pair $(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{I}(\Gamma))$ of correlation–automorphism groups of $\Gamma$ is isomorphic to $(G,H)$. Our construction proceeds in two main steps: first, we realize $(G,H)$ as the correlation and automorphism groups of an incidence system; then, we refine this system into a genuine incidence geometry preserving the same pair of groups. We also provide explicit examples, including a family of geometries realizing $(S_{n},A_{n})$ for all $n\geq 2$.
@@ 1. Introduction
In a recent paper [5], Leeman, Stokes, and Tranchida construed incident geometries for some groups. More precisely, they are interested in, for a given finite group $G$, finding an incident geometry $\Gamma$ such that the automorphism group of $G$ is isomorphic to the group of correlations of $\Gamma$ so that the inner automorphisms of $G$ correspond to the automorphisms of $\Gamma$. The authors hope that the combinatorial properties of such incident geometries are deeply connected to the algebraic structure of $G$ and its automorphisms. In [5], they constructed several incident geometries for classical families of groups and pave the way for the search of incident geometries for any group.
@@
Let $(G,H)$ be a pair of finite groups. There exists an incidence geometry $\Gamma$ such that $(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))\cong(G,H)$.
@@
Let $(G,H)$ be a pair of finite groups. If $G$ is trivial, the incidence geometry with one single element will suffice. Assume now that $G$ is non trivial. By Theorem 3.1, there exists an incidence system $\Gamma$ such that $(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))\cong(G,H)$ and such that all elements have degree at least 2, and $\Gamma$ has no flag of rank $3$ or more. Then by Theorem 4.1, there exists an incidence geometry $\Gamma^{\prime}$ such that $(\operatorname{Aut}(\Gamma^{\prime}),\operatorname{Aut}_{\text{I}}(\Gamma^{\prime}))\cong(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))\cong(G,H)$. ∎
The proof of Theorem A is inspired by the realizability of groups as groups of automorphisms of graphs [3]. It is based on the construction given in Theorem 3.1, which provides an incidence system with the necessary properties to apply Theorem 4.1. This latter theorem constructs, from an incident system satisfying some technical conditions on degrees and flags in the associated incident graph, an incident geometry with the same pair correlations-automorphisms.
@@ 2. Incidence graphs, incidence geometries and their automorphisms
3. Groups and normal subgroups as automorphisms of incidence systems
@@
Let $G$ be a finite non trivial group and $H\leq G$ be a normal subgroup, so they define a pair $(G,H)$. Then there exists an incidence system $\Gamma$ such that
@@
$V(\overline{\mathcal{G}})$ $=\{P_{i}\,:\,i=1,\ldots,\mathbf{g}\}$
$E(\overline{\mathcal{G}})$ $=\{(P_{i},P_{j})\,:\,i\neq j\}$
@@ 4. From incidence systems to incidence geometries
Below, we construct, for a given incidence system realizing a pair of groups, an incidence geometry realizing the same pair. Then main underlying idea is to replace every edge by a chamber.
@@
Let $\Gamma=(X,\ast,t,I)$ be an incidence system such that all elements have degree at least two and all flags have rank at most two, then there exists an incidence geometry $\Gamma^{\prime}=(X^{\prime},\ast^{\prime},t^{\prime},I)$ with $(\operatorname{Aut}(\Gamma^{\prime}),\operatorname{Aut}_{\text{I}}(\Gamma^{\prime}))\cong(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))$.
@@
Theorem 4.1constructs an incidence geometry from an incidence system that shares the same pair correlations-automorphisms, provided the original geometry contains only flags of rank at most two and elements of degree at least two. The next result provides a method to refine any incidence system so that Theorem 4.1 can be applied.
@@
For any incidence system $\Gamma$, there exists an incidence system $\Gamma^{\prime}$ with $(\operatorname{Aut}(\Gamma^{\prime}),\operatorname{Aut}_{\text{I}}(\Gamma^{\prime}))\cong(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))$ and such that all flags of $\Gamma^{\prime}$ have rank at most two and all elements of $\Gamma^{\prime}$ have degree at least two.
@@ 5. Examples
The following example demonstrates that, for certain pairs of groups $(G,H)$, ad-hoc constructions can yield incidence systems that are strictly smaller than those produced by Theorem 3.1, thereby offering a more efficient realization in specific cases.

### v2

# Every finite group is represented by a finite incidence geometry

@@ Abstract.
We investigate the relationship between finite groups and incidence geometries through their automorphism structures. Building upon classical results on the realizability of groups as automorphism groups of graphs, we develop a general framework to represent pairs of finite groups $(G,H)$, where $H\trianglelefteq G$, as pairs of correlation–automorphism groups of suitable incidence geometries. Specifically, we prove that for every such pair $(G,H)$, there exists a finite incidence geometry $\Gamma$ satisfying that the pair $(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{I}(\Gamma))$ of correlation–automorphism groups of $\Gamma$ is isomorphic to $(G,H)$. Our construction proceeds in two main steps: first, we realize $(G,H)$ as the correlation and automorphism groups of an incidence system; then, we refine this system into a genuine incidence geometry preserving the same pair of automorphisms groups. We also provide explicit examples, including a family of geometries realizing $(S_{n},A_{n})$ for all $n\geq 2$.
@@ 1. Introduction
In a recent paper [5], Leeman, Stokes, and Tranchida construed incident geometries for some finite groups. More precisely, they are interested in, for a given finite group $G$, finding an incident geometry $\Gamma$ such that the automorphism group of $G$ is isomorphic to the group of correlations of $\Gamma$ so that the inner automorphisms of $G$ correspond to the automorphisms of $\Gamma$. The authors hope that the combinatorial properties of such incident geometries are deeply connected to the algebraic structure of $G$ and its automorphisms. In [5], they constructed several incident geometries for classical families of groups and pave the way for the search of incident geometries for any group.
@@
Let $(G,H)$ be a pair of finite groups. There exists a finite incidence geometry $\Gamma$ such that $(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))\cong(G,H)$.
@@
Let $(G,H)$ be a pair of finite groups. If $G$ is trivial, the incidence geometry with one single element will suffice. Assume now that $G$ is non trivial. By Theorem 3.1, there exists a finite incidence system $\Gamma$ such that $(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))\cong(G,H)$ and such that all elements have degree at least 2, and $\Gamma$ has no flag of rank $3$ or more. Then by Theorem 4.1, there exists an incidence geometry $\Gamma^{\prime}$ such that $(\operatorname{Aut}(\Gamma^{\prime}),\operatorname{Aut}_{\text{I}}(\Gamma^{\prime}))\cong(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))\cong(G,H)$. ∎
The proof of Theorem A is inspired by the realizability of finite groups as groups of automorphisms of graphs [3]. It is based on the construction given in Theorem 3.1, which provides an incidence system with the necessary properties to apply Theorem 4.1. This latter theorem constructs, from an incident system satisfying some technical conditions on degrees and flags in the associated incident graph, an incident geometry with the same pair correlations-automorphisms.
@@ 2. Incidence graphs, incidence geometries and their automorphisms
3. Finite groups and normal subgroups as automorphisms of incidence systems
@@
Let $G$ be a finite non trivial group and $H\leq G$ be a normal subgroup, so they define a pair $(G,H)$. Then there exists a finite incidence system $\Gamma$ such that
@@
$V(\overrightarrow{\mathcal{G}})$ $=\{P_{i}\,:\,i=1,\ldots,\mathbf{g}\}$
$E(\overrightarrow{\mathcal{G}})$ $=\{(P_{i},P_{j})\,:\,i\neq j\}$
@@ 4. From incidence systems to incidence geometries
Below, we construct, for a given incidence system realizing a pair of finite groups, an incidence geometry realizing the same pair. Then main underlying idea is to replace every edge by a chamber.
@@
Let $\Gamma=(X,\ast,t,I)$ be a finite incidence system such that all elements have degree at least two and all flags have rank at most two, then there exists a finite incidence geometry $\Gamma^{\prime}=(X^{\prime},\ast^{\prime},t^{\prime},I)$ with $(\operatorname{Aut}(\Gamma^{\prime}),\operatorname{Aut}_{\text{I}}(\Gamma^{\prime}))\cong(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))$.
@@
Theorem 4.1constructs a finite incidence geometry from a finite incidence system that shares the same pair correlations-automorphisms, provided the original geometry contains only flags of rank at most two and elements of degree at least two. The next result provides a method to refine any incidence system so that Theorem 4.1 can be applied.
@@
For any finite incidence system $\Gamma$, there exists a finite incidence system $\Gamma^{\prime}$ with $(\operatorname{Aut}(\Gamma^{\prime}),\operatorname{Aut}_{\text{I}}(\Gamma^{\prime}))\cong(\operatorname{Aut}(\Gamma),\operatorname{Aut}_{\text{I}}(\Gamma))$ and such that all flags of $\Gamma^{\prime}$ have rank at most two and all elements of $\Gamma^{\prime}$ have degree at least two.
@@ 5. Examples
The following example demonstrates that, for certain pairs of finite groups $(G,H)$, ad-hoc constructions can yield incidence systems that are strictly smaller than those produced by Theorem 3.1, thereby offering a more efficient realization in specific cases.